# Uncovering Cultural Representation Disparities in Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) have demonstrated impressive capabilities across a range of tasks, yet concerns about their potential biases exist. This work investigates the extent to which prominent VLMs exhibit cultural biases by evaluating their performance on an image-based country identification task at a country level. Utilizing the geographically diverse Country211 dataset, we probe several large vision language models (VLMs) under various prompting strategies: open-ended questions, multiple-choice questions (MCQs) including challenging setups like multilingual and adversarial settings. Our analysis aims to uncover disparities in model accuracy across different countries and question formats, providing insights into how training data distribution and evaluation methodologies might influence cultural biases in VLMs. The findings highlight significant variations in performance, suggesting that while VLMs possess considerable visual understanding, they inherit biases from their pre-training data and scale that impact their ability to generalize uniformly across diverse global contexts.
@@ 1 Introduction
Figure 1: Average of country-wise cultural recognition accuracy by several VLMs
@@
Vision-Language Models (VLMs) have rapidly advanced, demonstrating exceptional capabilities in integrating visual and textual information for a wide array of tasks, from image captioning to visual question answering (Liu et al., 2024; Alayrac et al., 2022; Wang et al., 2024). These models are increasingly being deployed in diverse applications, impacting areas such as education, healthcare, and public services globally (Zhang et al., 2024). However, as their influence grows, so do concerns regarding their potential to perpetuate and even amplify societal biases present in their training data (Zhao et al., 2017; Zhou et al., 2022; Weng et al., 2024). Cultural and geographical biases are particularly deleterious, as they can lead to unequal performance and representation across different populations and regions of the world (AlKhamissi et al., 2024; Manvi et al., 2024). Defining "culture" is inherently complex, encompassing a broad spectrum of social norms, values, practices, languages, and historical contexts that shape the lived experiences of individuals and communities (Kroeber et al., 1985). Establishing culture in computational settings presents a persistent challenge due to its multifaceted and dynamic nature. Empirical studies employ tractable proxies such as demographic or geographic proxies to enable systematic analysis (Adilazuarda et al., 2024). While nation-level aggregation can mask sub-national heterogeneity, prior work in human–computer interaction and cultural analytics has demonstrated that country labels often serve as a practical proxy for coarse-grained cultural signals when large-scale analyses are required (Obradovich et al., 2022). To quantify cultural disparities in VLMs, we adopt image-based country identification as a concrete proxy task in which a model must infer an image’s country of origin solely from visual cues, while also providing a justification. Prior work has shown that geolocation tasks reveal representational imbalances in visual models, as performance often correlates with the prevalence of training data from different regions (Pouget et al., 2024). Using a dataset, with a balanced distribution of images from each country, we ensure that the observed accuracy gaps stem from model bias rather than dataset imbalance. Further, the complex nature of images used could measure VLMs’ ability to distinguish similar cultures/countries.
@@ 2 Related Works
Table 1: Comparison between current work and prior works
Recent work has increasingly explored the socio-cultural dimensions of Large Language Models (LLMs), including how they encode, express, and respond to culturally specific knowledge. Studies have examined value alignment (Choenni and Shutova, 2024), moral reasoning across languages (Agarwal et al., 2024), and cultural persona (AlKhamissi et al., 2024), while also uncovering strong Western biases in model outputs (Naous et al., 2024) which risk marginalizing cultural diversity if deployed in real world. There have also been efforts to address these concerns, like prompting based on ethnographic fieldwork (AlKhamissi et al., 2024) and fine-tuning culture-specific LLMs (Li et al., 2024a). Similar studies have been extended for Vision Language Models (VLMs) starting from (Liu et al., 2021) over cultural aspects, but in a weaker capacity (Nwatu et al., 2023) showed that CLIP (Radford et al., 2021) struggled in data for poor socio-economic groups worldwide in the Dollar Street dataset (Gaviria Rojas et al., 2022). State-of-the-art off-shelf VLMs score much higher on images depicting Western scenes than equivalent East-Asian scenes for every vision task, such as identification, question-answering, and art emotion classification (Ananthram et al., 2025). Similarly, (Liu et al., 2025) reveals that VLMs show stronger performance in Western concepts and weaker results in African and Asian contexts. These findings align with the fact that large pretraining corpora are dominated by high-resource languages and regions. Of the samples that can be geo-located in the OpenImages dataset Kuznetsova et al. (2020), 32% were from only the United States, and 60% came from only six Western countries (Shankar et al., 2017). Such imbalances translate into a “Western bias” in model behavior (de Vries et al., 2019).
@@
Impact of Evaluation approach :
Beyond dataset content, the format of evaluation also impacts bias measurement. In particular, many of the above benchmarks use multiple-choice or true/false questions, but this can mask a model’s true understanding. Since language choice can influence bias, benchmarks are often performed across multiple languages. (Romero et al., 2024) observed that LLaVA-1.5-7B showed a significant performance drop of 19.6% when prompted without multiple choices in the CVQA data set. Models also showed lower performance when prompted in native language of the image’s country of origin. However, (Ananthram et al., 2025) observed that prompting in a culturally closer language can reduce Western bias in some VLMs. It was also observed that people of different cultures are capable of differently capable of describing what they see in an image (van Miltenburg et al., 2017). We build on these insights by comparing open-ended vs. multiple-choice prompts (including “hard” questions with challenging distractors) and by evaluating in both English and native languages, to see how the prompting strategy affects cultural bias in VLMs.
@@ 3 Dataset Used
Existing benchmarks highlight cultural blind spots in VLMs, but they generally either cover fewer categories or countries or are restricted to specialized domains. Our work differs by using an image-based country-identification task over 211 countries, providing much broader geographic coverage and adversarial probing. Further, the datasets utilized in the prior works utilize a images that might be easier to classify, including but not limited to close up shots of food items, popular monuments being the primary object in an image etc… The dataset we utilized introduces a lot of noise and randomness in a majority of images as seen in Figure 2. For instance, the examples from Norway, India and Egypt might be easy to classify, but the examples from Afghanistan and Kuwait require grasping certain features and their associated knowledge i.e the headgear pattern of the Kuwait image and how it is different from other countries in the region. Th example from Afghanistan requires noticing the afghan flag, while the appearance of the person in the left may try to mislead the VLM due to an appearance of different ethnicity.
Figure 2: Few samples from the dataset with associated labels and annotated categories
@@ 4 Experiments
Image perturbations :The open-ended experiments were re-done with several adversarial changes such as :
  • Rotation by 90* clockwise
  • Rotation by 90* anti-clockwise
@@ 5 Results
The results obtained through each experimental setting over countries of each region can be seen in Table 2.
Region Open-Ended MCQs with MCQs with
Similar choices Random choices
North America 41.90 73.66 80.24
Central America 11.14 69.71 68.01
Caribbean 13.63 50.50 71.40
South America 20.36 70.94 68.67
Oceania 19.01 57.46 68.85
Western Europe 30.90 57.90 77.46
Northern Europe 25.33 60.61 79.41
Eastern Europe 26.57 53.39 75.88
Middle East 29.33 68.41 77.14
Central Asia 26.69 53.54 78.05
East Asia 43.58 71.55 83.80
Southeast Asia 41.74 67.47 81.70
South Asia 49.13 69.00 85.45
North Africa 31.86 54.28 78.91
Central Africa 11.84 56.95 68.22
Southern Africa 20.35 74.23 74.16
Overall 27.71 63.10 76.08
Table 2: Region wise accuracies through various experimental setting
@@ 5.1 Effect of Language of Inputs on Results
The average of country level accuracies compared to each language as input can be seen in Figure 3. The language used for inputs had a very little effect i.e <2% for all languages. But at a country level, most countries remained unaffected by language of the prompt to a large extent with change in accuracy less that 0.1%. The only cases with a noticeable change in accuracy are some but not all of the countries that speak the target language predominantly. This can be seen in Figure 4. For example, Changing the input language from English to Spanish improved accuracy for Spain but the change over Latin-American countries was negligible. Similarly, while switching to Portuguese had improved the accuracy for Brazil by a good amount, it lead to a drop in accuracy for Portugal. Overall, The input language did cause an improvement in some but not all countries primarily associated with the language used. The results also partially contradict prior findings that prompting in culturally similar languages reduces western bias (Ananthram et al., 2025).
Figure 3: Average of accuracies obtained using each VLM compared to the language of inputs : Similar Distractors (Right), Random Distractors (Left)
Figure 4: Changes in country-wise accuracies with changing input language compared to English
@@ 5.2 Effect of Image Perturbations on Results
Figure 15 and Figure 16 display the changes in accuracy observed due to gray-scaling and rotating the images compared to the original images. Input image perturbations can have a large impact on the country-level biases in VLMs. Further, It can be assumed that the VLMs tested are not robust enough towards image perturbations, with each country being effected at a different scale between each model/perturbation. The overall averages can also be seen in Figure 9, Figure 10 and Figure 11 respectively. The overall averages before and after these image perturbations across all of the VLMs indicate that gray-scaling leads to a minor drop in average accuracy while rotation of images lead to a major drop in average accuracy. But at a country level, only few countries’ accuracy were affected considerably by gray-scaling while drop in accuracies at a country level was more uniformly distributed due to rotations.
@@ 5.3 Effect of Input Variations on Confidence
Figure 5: Average Model Confidence : Normal(Left), Grayscale(Middle), Rotated(Right)
Category Description
Appearance (Attire) This includes attires of some people from the image, clothes being hanged in the background, etc.
Appearance (People) This includes appearance / visual perception of people’s ethnicity, presence of any celebrities, etc.
Architecture (Exterior) This includes building facades, monuments, bridges, outdoor structures, and any external architectural elements visible in the scene.
Architecture (Interior) This includes indoor environments such as rooms, corridors, lobbies, staircases, furniture layouts, and interior design details.
Landscape (Water) This includes bodies of water such as oceans, rivers, lakes, waterfalls, ponds, and any aquatic scenery.
Landscape (Air) This includes aerial or bird’s-eye views, landscapes captured from above (e.g., drone shots), clouds, sky scenes, and horizon vistas.
Landscape (Vegetation) This includes forests, grasslands, gardens, crops, shrubs, trees, foliage patterns, and any plant life or greenery visible in the scene.
Texts/Scripts/Posters This includes signs, banners, billboards, labels, handwritten or printed text, posters, and any other written or graphic messaging.
Patterns/Designs This includes decorative motifs, surface textures, fabric prints, wallpaper or tile patterns, abstract designs, and repetitive graphical elements.
Table 3: Categories used for categorizing the images and their descriptions
@@ 5.4 Image Feature categories VS accuracy
Figure 6: Accuracy of samples containing features of each category compared to image perturbations used : Originals (Left), Gray-scaling (Center), Rotation (Right)
Figure 7: Country wise response distribution compared to the model used.
Figure 8: An arrow map displaying what countries’ images were classified as India : External Architecture
@@ 5.5 Distribution of Predicted countries
The distribution of responses in an open ended approach by each model tested can be seen in Figure 7. The output distributions varied largely among each model tested, even among models of the same family (i.e between Gemma-3-27B, Gemma-3-12B and Aya-vision-32B, Aya-vision-8B). The results obtained contradict the usual assumption about western biases in generative models, and was observed over a few nations with likely high training data proportion.
@@ 5.6 Misclassification Analysis
The mapping of misclassification of samples was not limited to similar or neighboring nations. This can be observed in Figure 21 to Figure 36. These misclassifications varied by each individual feature and provide a better fine-grained insights of cultural biases. For instance, Apart from neighboring / similar countries most Images from Africa and rural regions of South America were classified as India. This can be seen in Figure 8 where out of the 600 images (100 * 6 models), roughly 80-120 belong to this category for most countries, while many countries had most of their misclassified as originating from India.

### v2

@@ Abstract
Vision-Language Models (VLMs) have demonstrated impressive capabilities across a range of tasks, yet concerns about their potential biases exist. This work investigates the extent to which prominent VLMs exhibit cultural biases by evaluating their performance on an image-based country identification task at a country level. Utilising the geographically diverse Country211 dataset, we probe several large vision language models (VLMs) under various prompting strategies: open-ended questions, multiple-choice questions (MCQs), including challenging setups like multilingual and adversarial settings. Our analysis aims to uncover disparities in model accuracy across different countries and question formats, providing insights into how training data distribution and evaluation methodologies might influence cultural biases in VLMs. The findings highlight significant variations in performance, suggesting that while VLMs possess considerable visual understanding, they inherit biases from their pre-training data and scale that impact their ability to generalize uniformly across diverse global contexts.
@@ 1 Introduction
Vision-Language Models (VLMs) have rapidly advanced, demonstrating exceptional capabilities in integrating visual and textual information for a wide array of tasks, from image captioning to visual question answering (Liu et al., 2024; Alayrac et al., 2022; Wang et al., 2024). These models are increasingly being deployed in diverse applications, impacting areas such as education, healthcare, and public services globally (Zhang et al., 2024).
Figure 1: Visualization of the average country-wise recognition accuracy across the VLMs studied in this paper. VLMs perform well at recognizing images from North American and Western European countries, but there are clear disparities in performance for African and Central American countries.
@@
However, as their influence grows, so do concerns regarding their potential to perpetuate and even amplify societal biases present in their training data (Zhao et al., 2017; Zhou et al., 2022; Weng et al., 2024). Cultural and geographical biases are of particular concern because they can lead to unequal performance and representation across different populations and regions of the world (AlKhamissi et al., 2024; Manvi et al., 2024). Defining "culture" is inherently complex, encompassing a broad spectrum of social norms, values, practices, languages, and historical contexts that shape the lived experiences of individuals and communities (Kroeber et al., 1985). Establishing culture in computational settings presents a persistent challenge due to its multifaceted and dynamic nature. Empirical studies employ tractable proxies such as demographic or geographic proxies to enable systematic analysis (Adilazuarda et al., 2024; Yadav et al., 2025). While nation-level aggregation can mask sub-national heterogeneity, prior work in human–computer interaction and cultural analytics has demonstrated that country labels often serve as a practical proxy for coarse-grained cultural signals when large-scale analyses are required (Obradovich et al., 2022).
In order to quantify cultural disparities in VLMs, we adopt image-based country identification as a concrete proxy task in which a model must infer an image’s country of origin solely from visual cues, while also providing a justification. Prior work has shown that geolocation tasks reveal representational imbalances in visual models, as performance often correlates with the prevalence of training data from different regions (Pouget et al., 2024). Using a dataset, with a balanced distribution of images from each country, we ensure that the observed accuracy gaps stem from model bias rather than dataset imbalance. Further, the complex nature of images used could measure VLMs’ ability to distinguish similar cultures/countries.
The main contributions of this paper are:
  1. 1.
    We introduce a scalable framework to evaluate cultural biases in VLMs using an image-based country identification task over 211 countries, leveraging the geographically diverse and balanced Country211 dataset.
  2. 2.
    We systematically probe VLMs under varied settings—open-ended and multiple-choice questions (MCQs) with both random and culturally similar distractors—alongside multilingual prompts in five languages, to capture nuanced cultural and linguistic disparities.
  3. 3.
    We examine model robustness to image perturbations and analyse performance across nine image categories (e.g. architecture, landscape, food etc), revealing the influence of image content on cultural bias.
  4. 4.
    Our findings show that VLM biases do not consistently favour Western countries; instead, biases often reflect overrepresentation of certain popular countries (e.g., India, USA) in the training data, suggesting a more complex bias landscape.
@@ 2 Related Works
Table 1: Overview of prior datasets used in cultural recognition experiments.
Recent work has increasingly explored the socio-cultural dimensions of Large Language Models (LLMs), including how they encode, express, and respond to culturally specific knowledge. Studies have examined value alignment (Choenni and Shutova, 2024), moral reasoning across languages (Agarwal et al., 2024), and cultural persona (AlKhamissi et al., 2024), while also uncovering strong Western biases in model outputs (Naous et al., 2024) which risk marginalizing cultural diversity if deployed in real world. There have also been efforts to address these concerns, like prompting based on ethnographic fieldwork (AlKhamissi et al., 2024) and fine-tuning culture-specific LLMs (Li et al., 2024a). Similar studies have been extended for Vision Language Models (VLMs) starting from (Liu et al., 2021) over cultural aspects, but in a weaker capacity (Nwatu et al., 2023) showed that CLIP (Radford et al., 2021) struggled in data for poor socio-economic groups worldwide in the Dollar Street dataset (Gaviria Rojas et al., 2022). State-of-the-art off-shelf VLMs score much higher on images depicting Western scenes than equivalent East-Asian scenes for every vision task, such as identification, question-answering, and art emotion classification (Ananthram et al., 2025). Similarly, (Liu et al., 2025; Yadav et al., 2025) reveals that VLMs show stronger performance in Western concepts and weaker results in African and Asian contexts. These findings align with the fact that large pretraining corpora are dominated by high-resource languages and regions. Of the samples that can be geo-located in the OpenImages dataset Kuznetsova et al. (2020), 32% were from only the United States, and 60% came from only six Western countries (Shankar et al., 2017). Such imbalances translate into a “Western bias” in model behavior (de Vries et al., 2019).
@@
Figure 2: Examples of the Country211 dataset, alongside automatically-predicted categories for each image, showcasing the visual diversity of the examples to be classified.
@@
Impact of Evaluation:
The format of evaluation also impacts bias measurement. Many of the above benchmarks use multiple-choice or binary questions, which can mask a model’s true understanding. Since language choice can influence bias, benchmarks are often performed across multiple languages. (Romero et al., 2024) showed that the performance of LLaVA-1.5-7B dropped by 19.6% when prompted without multiple choices for CVQA. Models also showed lower performance when prompted in native language of the image’s country of origin. However, (Ananthram et al., 2025) observed that prompting in a culturally closer language can reduce Western bias in some VLMs. It was also observed that people of different cultures are capable of differently capable of describing what they see in an image (van Miltenburg et al., 2017). We build on these insights by comparing open-ended vs. multiple-choice prompts (including “hard” questions with challenging distractors) and by evaluating in both English and native languages, to see how the prompting strategy affects cultural bias in VLMs.
@@ 3 Dataset Used
Existing benchmarks highlight cultural blind spots in VLMs, but they generally either cover fewer categories or countries or are restricted to specialized domains. Our work differs by using an image-based country-identification task over 211 countries, providing much broader geographic coverage and adversarial probing. Further, the datasets utilized in the prior works utilize a images that might be easier to classify, including but not limited to close up shots of food items, popular monuments being the primary object in an image etc. The dataset we utilized introduces a lot of noise and randomness in a majority of images as seen in Figure 2. For instance, the examples from Norway, India and Egypt might be easy to classify, but the examples from Afghanistan and Kuwait require grasping certain features and their associated knowledge i.e the headgear pattern of the Kuwait image and how it is different from other countries in the region. Th example from Afghanistan requires noticing the afghan flag, while the appearance of the person in the left may try to mislead the VLM due to an appearance of different ethnicity.
Figure 3: Model-wise averaged accuracy when varying the prompt language or selection of MCQA alternatives (left: random; right: similar). Performance is consistent across conditions.
@@ 4 Experiments
Image perturbations:The open-ended experiments were re-done with following adversarial changes:
  1. 1.
    Rotation by 90 clockwise,
  2. 2.
    Rotation by 90 anti-clockwise
  3. 3.
@@ 5 Results
The results for experimental setting over countries of each region can be seen in Table 2.
MCQA
Region Open-Ended Similar Random
North America 41.9 73.7 80.2
Central America 11.1 69.7 68.0
Caribbean 13.6 50.5 71.4
South America 20.4 70.9 68.7
Oceania 19.0 57.5 68.9
Western Europe 30.9 57.9 77.5
Northern Europe 25.3 60.6 79.4
Eastern Europe 26.6 53.4 75.9
Middle East 29.3 68.4 77.1
Central Asia 26.7 53.5 78.1
East Asia 43.6 71.6 83.8
Southeast Asia 41.7 67.5 81.7
South Asia 49.1 69.0 85.5
North Africa 31.9 54.3 78.9
Central Africa 11.8 57.0 68.2
Southern Africa 20.4 74.2 74.2
Overall 27.7 63.1 76.1
Table 2: Region-wise averaged accuracy across models. There are consistent disparities in performance across different regions, regardless of the prompting method.
@@ 5.1 Effect of Language of Inputs on Results
The average of country level accuracies compared to each language as input can be seen in Figure 3. The language used for inputs had a very little effect i.e <2% for all languages. But at a country level, most countries remained unaffected by language of the prompt to a large extent with change in accuracy less that 0.1%. The only cases with a noticeable change in accuracy are some but not all of the countries that speak the target language predominantly. For example, Changing the input language from English to Spanish improved accuracy for Spain but the change over Latin-American countries was negligible. Similarly, while switching to Portuguese had improved the accuracy for Brazil, it lead to a drop in accuracy for Portugal. Overall, the input language improves performance for some countries primarily associated with the language used. The results also partially contradict prior findings that prompting in culturally similar languages reduces western bias (Ananthram et al., 2025).
Figure 4: Model-wise averaged accuracy across the nine image categories, as a function of the image perturbations. There is a clear trend of models performing better with the original images (left), compared to the grayscale images (middle) , or rotated images (right).
@@ 5.2 Effect of Image Perturbations on Results
Figure 4 and Figure 5 display the changes in accuracy observed due to gray-scaling and rotating the images compared to the original images. Input image perturbations can have a large impact on the country-level biases in VLMs. Further, It can be assumed that the VLMs tested are not robust enough towards image perturbations, with each country being effected at a different scale between each model/perturbation. The overall averages can also be seen in Figure 8, Figure 9 and Figure 10 respectively.
Figure 18 shows how perturbations affect model performance across different semantic image categories. For all nine categories, models perform best on original (unaltered) images, with decreasing accuracy for gray-scaled and worse for rotated versions. Categories like exterior architecture, text/scripts/posters, and attire/patterns are especially impacted by perturbations. We hypothesize that it is likely because they contain fine-grained, orientation-sensitive, or highly color-dependent details.
We also look at geographical disparities of these changes in orientation in Figure 14 and  Figure 15. We observe the disparity in model robustness also emerges clearly. For example, models such as Aya Vision 32B, GPT4o-mini and Gemini 3 12B show very different sensitivity across both a) perturbations and b) regions which were affected. We hypothesise that architectural and training differences might be influencing how models process image orientation and color. While gray-scaling may reduce performance due to the loss of visual detail or color-dependent cues, rotation disrupts spatial reasoning and object orientation, which are critical for geographic or cultural recognition.
These findings highlight the importance of evaluating model performance under realistic image distortions, especially for applications where images may not be clean or consistently formatted as image characteristics can vary widely.
Figure 5: Average model confidence, given the original images (left), grayscale images (middle), and rotated images (right). GPT4o, GeminiFlash, and Gemma27B are most sensitive to image perturbations.
@@ 5.3 Effect of Input Variations on Confidence
Figure 6: Country-wise response distribution in the open-ended prompt format. There is a consistent trend of models predicting USA, but otherwise, no clear bias towards predicting Western countries.
@@ 5.5 Distribution of Predicted countries
The distribution of responses in an open ended approach can be seen in Figure 6. The output distributions varied largely among models, even those within the same family (i.e between Gemma-3-27B, Gemma-3-12B and Aya-vision-32B, Aya-vision-8B). The results obtained contradict the usual assumption about western biases in generative models, and was observed over a few nations with likely high training data proportion.
Notably, all models consistently overpredict certain countries—particularly the USA, India, and Brazil—regardless of actual ground truth. We hypothize that these countries are likely overrepresented in the models’ pretraining data or benefit from more visually distinctive cues. Biases seem to cluster around a few highly represented or visually salient countries rather than reflecting broader geopolitical landscape.
These results show that model predictions are likely highly influenced by data availability and image characteristics rather than a generic global bias. It also underscores the need for better interpretabilty regarding the geographic composition of VLM training datasets to fully understand such biases.
@@ 5.6 Misclassification Analysis
Figure 7: Mis-classification map for North African countries. There is a clear trend of models predicting USA, India, Australia, or geographically close countries in Europe and the Middle East.
The mapping of misclassification of samples was not limited to similar or neighboring nations. This can be observed in Figure 20 to Figure 34. These misclassifications varied by each individual feature and provide a better fine-grained insights of cultural biases. For instance, Apart from neighboring / similar countries, most images from Africa and rural regions of South America were classified as India. A specific example is shown in Figure 7 where out of the 600 images (100 * 6 models), roughly 80-120 belong to this category for most countries, while many countries had most of their misclassified as originating from India.
6 Discussion
Our study presents a comprehensive analysis of cultural biases in Vision-Language Models (VLMs) using a geographically balanced dataset across 211 countries. The key difference between the samples in our datasets and those from the other works is the complexity of images provided. In most of the images, we used it is difficult to understand and predict the country of origin without a strong vision ability. For example, the Afghanistan example from Figure 2 is difficult to classify without spotting the hint of Afghan flag from one of the people from the image.
We evaluated popular models across multiple prompting strategies, e.g. open-ended, multiple-choice (random and similar distractors), and multilingual settings. Open-ended formats showed the greatest disparity in country-level accuracy, particularly in underrepresented regions such as Central Africa and parts of South America.The use of culturally similar distractors proved to be the most effective in revealing fine-grained errors, highlighting limitations in models’ cultural discrimination abilities.
We further assessed the models’ robustness to image perturbations like gray-scaling and rotation. While gray-scaling affected only a few specific countries, rotation led to a broad and uniform drop in performance, confirming that VLMs rely heavily on image orientation. We further observed that performance also varied by semantic image content—categories like architecture, textual cues, and attire were more predictive of cultural origin, especially in unaltered images.
Language variation in prompts had minimal impact on average accuracy, though countries closely tied to the input language (e.g., Spain with Spanish, Brazil with Portuguese) showed slight gains. However, this trend was inconsistent and did not generalize across all culturally linked regions.
Finally, our misclassification analysis shows that models frequently confuse images from low-resource or visually ambiguous countries with a few dominant nations, reinforcing the role of training data bias. These findings emphasize the need for transparent reporting of dataset composition in VLM development and call for more robust, culturally diverse benchmarks to ensure equitable global performance